Operations Report – April 21st

Crew 297 Operations Report 21-04-2024

SOL: 7

Name of person filing report: Matt Lynch (Engineer)

Non-nominal systems: Suit 2 (stuck valve), Suit 9 (broken visor),

Notes on non-nominal systems: The issues with Suit 2 and 9 were already noted before our mission. Suit 11’s battery has now been replaced


Opportunity rover used:

Hours: 182.7

Beginning charge: 100%

Ending charge: 43%

Currently charging: No

General notes on rovers: All rovers key locks were tighten today on EVA #10

Summary of Hab operations:

WATER USE: 25.4 gallons

Water (static tank): 258.8 gallons

Static tank pipe heater (on or off): off

Static tank heater (On or off): off

Toilet tank emptied: yes

Summary of internet: Looking up recipes, downloading entertainment, report formatting methods, catching up on Earthly news

Summary of suits and radios:

Radios 2,3,4,5, and 6 were used.

EVA Suits 1,4,6 and 7 were used.

Summary of GreenHab operations:

WATER USE: 4 gallons

Heater: Yes

Supplemental light: N/A

Harvest: 106 g of cherry tomatoes

Summary of ScienceDome operations: NA

Summary of RAM operations: (Tools used, work done): Tools from the RAM were used in suit repair today. EVA #10 used the RAM to perform their robotic studies in the area right in front of the Hab.

Summary of any observatory issues: NA

Copyright © The Mars Society. All rights reserved. | Main Site